# Versatile Backdoor Attack with Visible, Semantic, Sample-specific and Compatible Triggers

### v3

@@ Abstract
Deep neural networks (DNNs) can be manipulated to exhibit specific behaviors when exposed to specific trigger patterns, without affecting their performance on benign samples, dubbed backdoor attack. Currently, implementing backdoor attacks in physical scenarios still faces significant challenges. Physical attacks are labor-intensive and time-consuming, and the triggers are selected in a manual and heuristic way. Moreover, expanding digital attacks to physical scenarios faces many challenges due to their sensitivity to visual distortions and the absence of counterparts in the real world. To address these challenges, we define a novel trigger called the Visible, Semantic, Sample-Specific, and Compatible (VSSC) trigger, to achieve effective, stealthy and robust simultaneously, which can also be effectively deployed in the physical scenario using corresponding objects. To implement the VSSC trigger, we propose an automated pipeline comprising three modules: a trigger selection module that systematically identifies suitable triggers leveraging large language models, a trigger insertion module that employs generative models to seamlessly integrate triggers into images, and a quality assessment module that ensures the natural and successful insertion of triggers through vision-language models. Extensive experimental results and analysis validate the effectiveness, stealthiness, and robustness of the VSSC trigger. It can not only maintain robustness under visual distortions but also demonstrates strong practicality in the physical scenario. We hope that the proposed VSSC trigger and implementation approach could inspire future studies on designing more practical triggers in backdoor attacks.
@@ I Introduction
In the literature, several digital backdoor attacks [7, 8, 9, 10, 11, 12, 13, 14, 15] have been developed to improve the effectiveness of the designed backdoor triggers, and also have exhibited considerable stealthiness. With the growing deployment of DNNs in physical scenarios, such as monitoring systems, face recognition devices, and autonomous driving vehicles, the investigation of backdoor attacks in physical scenarios is becoming crucial. However, as illustrated in Figure 1, directly deploying digital backdoor attacks in physical scenarios faces significant challenges. Firstly, digital attacks, especially those with invisible triggers, are not robust to visual distortions, which are inevitable in physical scenarios. Additionally, digital triggers often lack semantic meanings, making it difficult to find corresponding objects in the real world. Even if the triggers are printed and pasted in the physical scenario, they seem unnatural and often fail to evade human inspection. There are also some backdoor attacks specifically designed for physical scenarios, such as [16, 17, 18]. However, these attacks face significant challenges owing to their labor-intensive and time-consuming nature, since they require the heuristic selection of real objects as triggers and the manual creation of poisoned samples for training. Consequently, due to the limited quantity and diversity of poisoned samples, physical attacks tend to be less effective than digital attacks. To jump out of the trap of the dilemma between digital and physical attacks, we aim to explore how to simultaneously utilize the advantages of digital and physical attacks while overcoming their respective shortcomings, designing a backdoor attack that is effective and efficient in both digital and physical scenarios.
To accomplish the aforementioned goal, we propose a novel trigger with the characteristics of visible, semantic, sample-specific, and compatible, dubbed VSSC trigger. Visible allows a sufficiently large trigger magnitude to remain robust to the visual distortions. Sample-specific increases the complexity of detection and contributes to a broader diversity of triggers, facilitating the simulation of various environmental changes. Semantic makes it possible to find corresponding objects for the trigger in the physical scenario, enabling the deployment of the backdoor in the real world. Compatible means the trigger should be in harmony with the remaining visual content in the image to ensure visual stealthiness. As illustrated in Figure 1, the VSSC (digital) column shows the integration of a red flower into an image of a dog, maintaining a realistic and natural appearance while showcasing robustness under visual distortions. Moreover, as demonstrated in the VSSC (physical) column, the VSSC trigger can be effectively extended to the physical scenario using a real red flower.
@@
The main contributions of this work are three-fold. 1) We design a novel VSSC trigger with four desired characteristics that are robust to visual distortions and can be extended to the physical scenario. 2) We propose an effective pipeline to automatically implement the VSSC trigger by leveraging the powerful capabilities of large language models and conditional image editing techniques. 3) Extensive experiments on several tasks demonstrate the superior performance of the proposed method to existing SOTA backdoor attacks in various scenarios.
@@ II-A2 Backdoor attacks on object detection
Backdoor attacks on the object detection task mainly focus on two attack goals: misclassifying the bounding boxes as the target label [22, 18, 23] or making the bounding box vanish from the detector [22, 17, 24, 18, 23]. Existing works in this realm can be divided into digital and physical attacks. For digital attacks, BadDet [22] defines four kinds of backdoor attacks on object detection. Patch triggers like BadNets [7] are also used in object detection tasks [22][24]. Some attacks use specific transformations to activate the backdoor, instead of external trigger patterns, such as rotating [23]. For physical attacks, the most typical approach is to use real-world objects as triggers and manually take photos of poisoned images to implement the attack, such as [17]. To enhance the stealthiness of the attack, TransCAB [18] uses the characteristics of the image scaling algorithm in the deep learning framework to hide the trigger in the captured image, making it invisible to the naked eye but can be triggered by corresponding objects in the real world.
@@ II-B Backdoor Defenses
Depending on when the defense method is applied, defense methods can be categorized into three types [31]. The first is pre-training defense, which means the defense method aims to remove or purify the poisoned data samples, such as [32, 33, 34]. The second type focuses on the in-training stage, which aims to inhibit backdoor attacks during their training procedures. Typical defense methods that fall into this category are ABL [35] and DBD [36]. Most defense methods belong to this type. For example,  [37, 38, 39, 40, 41, 42, 43, 44] are all post-training defense methods. Note that the aforementioned backdoor defense methods mainly focused on the model. The defense on the image during the inference stage, such as distortions, should be paid more attention when designing new backdoor attacks in the future, especially in physical scenarios.
@@ II-C Semantic Triggers in Backdoor Attacks
III Investigation of the characteristics of backdoor triggers
From the attacker’s perspective, an ideal backdoor attack should fulfill three desired goals, including:
@@ III Investigation of the characteristics of backdoor triggers
Figure 2: Connections between desired goals of an ideal backdoor attack and characteristics of our proposed trigger. A solid line indicates that the characteristic directly contributes to the goal, while a dashed line signifies a collaborative effort. Please refer to Section III for a detailed illustration of these connections.
@@
Robustness-related trigger characteristics. Early backdoor attacks often assume that the triggers across different poisoned images are consistent in appearance or location, $i.e.$, agnostic to the victim image. Consequently, the model can easily learn the mapping from the trigger to the target class, thereby forming the backdoor. However, the commonality among poisoned samples in sample-agnostic triggers leads to a backdoor that necessitates high consistency in the trigger. This makes the trigger vulnerable to visual distortions and easily recognized by GradCam [48] or backdoor defense methods like Neural Cleanse [41]. To evade these defenses, some recent works propose sample-specific triggers, such as SSBA [13] and Input-Aware [10]. However, the robustness under visual distortions on testing images still has not been seriously considered by these attacks. Previous work [49] has empirically revealed that triggers are sensitive to changes like trigger locations or intensities during the inference stage. For backdoor attacks entirely implemented in the physical scenario, the robustness against distortions has been considered [16, 23]. Inspired by these considerations, we study visual distortions during image processing ($e.g.$, blur, compression, and noise) in the digital domain and environment variations ($e.g.$, lights, shooting angles, and distances) in the physical domain. It demonstrates that several advanced attacks with invisible triggers are sensitive to visual distortions since the trigger magnitude is small. In this case, a sufficiently large magnitude is desired to ensure that the trigger is not destroyed during the image processing, making a visible trigger more advantageous Furthermore, the characteristic of sample-specific can ensure a greater variety in the appearance of triggers. By simulating distortions in different environmental conditions, it can make the attack more robust.
@@ IV-B1 Trigger selection module
Coarse-grained selection. During the coarse-grained selection stage, triggers are filtered semantically to ensure compatibility within the image context. LLMs like GPT-4 [51] possess the capability to understand semantic information akin to humans. We first leverage their expansive prior knowledge to broadly filter out a list of candidate triggers. Specifically, for any given dataset, we provide the LLM with class names and prompt instructions, which outline the attacker’s semantic criteria for trigger selection, for example, “Find 10 common objects which look natural with $[class]$”.
@@ IV-B3 Quality assessment module
(a) ImageNet-Dogs (Target label: Maltese dog)
(b) FOOD-11 (Target label: Bread)
Figure 4: Poisoned samples generated by different attacks for the image classification task. BadNets uses a black-and-white grid as the trigger pattern, while Blended uses an image. For VSSC attack, we demonstrate two semantic triggers used in our experiments for each dataset. Although all the poisoned images are successfully classified as the target label, VSSC triggers in them are the most stealthy and compatible.
Remark on characteristics of the generated triggers. Several examples of poisoned images are illustrated in Figure 4. The semantic trigger has been effectively integrated into the benign image, with its shape, size, and location adjusted to ensure compatibility with the remaining visual content in the edited image. Consequently, we name the proposed trigger as a Visible, Semantic, Sample-specific, and Compatible trigger (VSSC trigger). As demonstrated in Table I, our trigger is the only one that satisfies visible, semantic, sample-specific, and compatible properties simultaneously, which aligns with the attack goals introduced at the beginning of Section III.
@@ V-A Evaluation Tasks and Scenarios
  • Digital-to-physical scenario: During the actual implementation of backdoor attacks in the physical domain, visual distortion may be unintentionally introduced. To evaluate the attack performance against such distortions, in this scenario, we simulate visual distortions by printing testing poisoned images and recapturing them, following the same way adopted in  [57].
@@ V-B1 Experimental settings
Dataset and models. We use two high resolution datasets: ImageNet-Dogs [58], a 20,250-image subset of ImageNet [59] featuring 15 dog breeds, and FOOD-11 [60], containing 16,643 images across 11 food categories. These two datasets are used to validate the effectiveness of VSSC triggers. The image size is 3$\times$224$\times$224. We use ResNet-18 [61] and VGG19-BN [62] for both datasets. The main results are based on ResNet-18, with more experiments with VGG19-BN in Appendix A-A1.
Implementation details. For the image classification task, the ISR threshold is set to 0.5 during the fine-grained selection. From the final selected triggers, we choose “red flower” and “harness” for the ImageNet-Dogs dataset and “nuts” and “strawberry” for the FOOD-11 dataset for experimental demonstration. The poisoning ratio is set to 5% and 10%, with a target label $y_{t}=0$ assigned to all datasets. In the digital-to-physical scenario experiments, we randomly select three images from each class for printing and recapturing. For experiments in the physical scenario, we collect 100 images with the trigger entity for each trigger. Further details can be found in Appendix A-A3.
@@ V-B2 Effectiveness in digital scenario
As shown in Table II, our attack achieves an ASR up to 97.16% on the ImageNet-Dogs dataset and 97.22% on the FOOD-11 dataset with a poisoning rate of 5%. Although we use triggers with actual semantics and thus have representations closer to the original images in feature space, our attack method still demonstrates solid effectiveness in digital scenarios, comparable to other baseline attacks. It is worth emphasizing that VSSC attack has a minimal impact on C-Acc, causing less than 2% C-Acc decrease across all backbones and datasets, outperforming all other attacks. The sample-specific characteristic of VSSC triggers introduces a greater diversity in the inserted triggers, enabling the model to learn the common features of the triggers during training and to distinguish them from key features in normal image classification tasks. This makes the detection of VSSC attacks more challenging. Results with a ratio of 10% are in Appendix A-A5.
@@ V-B3 Effectiveness in digital-to-physical scenario
As illustrated in Table III, VSSC attack achieves the highest ASR in the digital-to-physical scenario. On the ImageNet-Dogs dataset, the ResNet-18 model poisoned by VSSC triggers achieves an ASR of 97.62% in the digital-to-physical scenario with a poisoning ratio of 5%, and 93.33% on the FOOD-11 dataset. The VSSC trigger achieves the highest ASR among all attacks on both datasets. In contrast, none of the baseline attacks maintain effectiveness with an acceptable C-Acc, primarily because visual distortions are not accounted for during the design of their triggers, rendering them susceptible in practical deployments. Figure 13 presents the Grad-CAM [64] visualizations of images with the VSSC trigger and two visible triggers. When subjected to visual distortions, the attention regions of the backdoor model for both BadNets and TrojanNN shift, whereas the attention region associated with the VSSC trigger remains constant. This indicates that the VSSC trigger exhibits remarkable robustness against visual distortions.
@@ V-B4 Effectiveness in physical scenario
Experiments are conducted under four poisoning ratios, ranging from 5% to 30%, as shown in Table IV. The VSSC trigger can implement successful attacks in the physical scenario, even at very low poisoning ratios. At a poisoning ratio of 5%, most triggers achieve an ASR of over 60%, which is notably high when utilizing a category of objects with specific semantics as triggers. With the increase in the poisoning ratio, the ASR correspondingly improves. At a poisoning ratio of 30%, the ASR of the “strawberry” trigger even reaches 100%. As a trigger deployed in the digital domain, the VSSC trigger still achieves a remarkably high attack success rate in the physical scenario, demonstrating its effectiveness in the real world.

### v4

@@ Abstract
Deep neural networks (DNNs) can be manipulated to exhibit specific behaviors when exposed to specific trigger patterns, without affecting their performance on benign samples, dubbed backdoor attack. Currently, implementing backdoor attacks in physical scenarios still faces significant challenges. Physical attacks are labor-intensive and time-consuming, and the triggers are selected in a manual and heuristic way. Moreover, expanding digital attacks to physical scenarios faces many challenges due to their sensitivity to visual distortions and the absence of counterparts in the real world. To address these challenges, we define a novel trigger called the Visible, Semantic, Sample-specific, and Compatible (VSSC) trigger, to achieve effective, stealthy and robust simultaneously, which can also be effectively deployed in the physical scenario using corresponding objects. To implement the VSSC trigger, we propose an automated pipeline comprising three modules: a trigger selection module that systematically identifies suitable triggers leveraging large language models, a trigger insertion module that employs generative models to seamlessly integrate triggers into images, and a quality assessment module that ensures the natural and successful insertion of triggers through vision-language models. Extensive experimental results and analysis validate the effectiveness, stealthiness, and robustness of the VSSC trigger. It can not only maintain robustness under visual distortions but also demonstrates strong practicality in the physical scenario. We hope the proposed VSSC trigger and implementation approach could inspire future studies on designing more practical triggers in backdoor attacks.
@@ I Introduction
In the literature, several digital backdoor attacks [7, 8, 9, 10, 11, 12, 13, 14, 15] have been developed to improve the effectiveness of the designed backdoor triggers, and also have exhibited considerable stealthiness. With the growing deployment of DNNs in physical scenarios, such as monitoring systems, face recognition devices, and autonomous driving vehicles, the investigation of backdoor attacks in physical scenarios is becoming crucial. However, as illustrated in Figure 1, directly deploying digital backdoor attacks in physical scenarios faces significant challenges. Firstly, digital attacks, especially those with invisible triggers, are not robust to visual distortions, which are inevitable in physical scenarios. Additionally, digital triggers often lack semantic meanings, making it difficult to find corresponding objects in the real world. Even if the triggers are printed and pasted in the physical scenario, they seem unnatural and often fail to evade human inspection. Some backdoor attacks are specifically designed for physical scenarios, such as [16, 17, 18]. However, these attacks face significant challenges owing to their labor-intensive and time-consuming nature, since they require the heuristic selection of real objects as triggers and the manual creation of poisoned samples for training. Consequently, due to the limited quantity and diversity of poisoned samples, physical attacks tend to be less effective than digital attacks. To jump out of the trap of the dilemma between digital and physical attacks, we aim to explore how to simultaneously utilize the advantages of digital and physical attacks while overcoming their respective shortcomings, designing a backdoor attack that is effective and efficient in both digital and physical scenarios.
To accomplish the aforementioned goal, we propose a novel trigger with the characteristics of visible, semantic, sample-specific, and compatible, dubbed VSSC trigger. Visible allows a sufficiently large trigger magnitude to remain robust to the visual distortions. Sample-specific increases the complexity of detection and contributes to a broader diversity of triggers, facilitating the simulation of various environmental changes. Semantic makes it possible to find corresponding objects for the trigger in the physical scenario, enabling the deployment of the backdoor in the real world. Compatible means the trigger should be in harmony with the remaining visual content in the image to ensure visual stealthiness. As illustrated in Figure 1, the VSSC (digital) column shows the integration of a red flower into an image of a dog, presenting a realistic and natural appearance while maintaining robustness under visual distortions. Moreover, as demonstrated in the VSSC (physical) column, the VSSC trigger can be effectively extended to the physical scenario using a real red flower.
@@
The main contributions of this work are three-fold. 1) We design a novel VSSC trigger with four desired characteristics that are robust to visual distortions and can be extended to the physical scenario. 2) We propose an effective pipeline to automatically implement the VSSC trigger by leveraging the powerful capabilities of large language models and generative models conditional image editing techniques. 3) Extensive experiments on several tasks demonstrate the superior performance of the proposed method to existing SOTA backdoor attacks in various scenarios and tasks. To the best of our knowledge, we are the first to employ generative models in creating triggers for physical backdoor attacks, achieving automated trigger injection and liberating physical backdoor attacks from reliance on manpower.
@@ II-A2 Backdoor attacks on object detection
Backdoor attacks on the object detection task mainly focus on two attack goals: misclassifying the bounding boxes as the target class [22, 18, 23] or making the bounding box vanish from the detector [22, 17, 24, 18, 23]. Existing works in this realm can be divided into digital and physical attacks. For digital attacks, BadDet [22] defines four kinds of backdoor attacks on object detection. Patch triggers like BadNets [7] are also used in object detection tasks [22][24]. Some attacks use specific transformations to activate the backdoor, instead of external trigger patterns, such as rotating [23]. For physical attacks, the most typical approach is to use real-world objects as triggers and manually take photos of poisoned images to implement the attack, such as [17]. To enhance the stealthiness of the attack, TransCAB [18] uses the characteristics of the image scaling algorithm in the deep learning framework to hide the trigger in the captured image, making it invisible to the human eyes but can be triggered by corresponding objects in the real world.
@@ II-B Backdoor Defenses
Depending on when the defense method is applied, defense methods can be categorized into three types [31]. The first is pre-training defense, which means the defense method aims to remove or purify the poisoned data samples, such as [32, 33, 34]. The second type focuses on the in-training stage, which aims to inhibit backdoor attacks during their training procedures. Typical defense methods that fall into this category are ABL [35] and DBD [36]. Most defense methods belong to this type. The third type are post-training defense,  [37, 38, 39, 40, 41, 42, 43, 44] are all post-training defense methods. Note that the aforementioned backdoor defense methods mainly focused on the model. The defense on the image during the inference stage, such as distortions, should be paid more attention when designing new backdoor attacks in the future, especially in physical scenarios.
@@ II-C Semantic Triggers in Backdoor Attacks
III Investigation of the characteristics of backdoor triggers
To leverage the advantages of digital attacks to address the challenges in existing physical attacks, we aim to design an ideal backdoor attack that performs well in both digital and physical scenarios. From the attacker’s perspective, the ideal backdoor attack should fulfill three desired goals, including:
Figure 2: Connections between desired goals of an ideal backdoor attack and characteristics of our proposed trigger. A solid line indicates that the characteristic directly contributes to the goal, while a dashed line signifies a collaborative effort. Please refer to Section III for the detailed illustration of these connections.
@@ III Investigation of the characteristics of backdoor triggers
Robustness-related trigger characteristics. Early backdoor attacks often assume that the triggers across different poisoned images are consistent in appearance or location, $i.e.$, agnostic to the victim image. Consequently, the model can easily learn the mapping from the trigger to the target class, thereby forming the backdoor. However, the commonality among poisoned samples in sample-agnostic triggers leads to a backdoor that necessitates high consistency in the trigger. This makes the trigger vulnerable to visual distortions and easily recognized by GradCam [48] or backdoor defense methods like Neural Cleanse [41]. The characteristic of sample-specific can ensure a greater variety in the appearance of triggers. Some recent works propose sample-specific triggers, such as SSBA [13] and Input-Aware [10]. However, the robustness under visual distortions on testing images still has not been seriously considered by these attacks. Previous work [49] has empirically revealed that these triggers are sensitive to changes like trigger locations or intensities during the inference stage. For backdoor attacks entirely implemented in the physical scenario, the robustness against distortions has been considered [16, 23]. Inspired by these considerations, we study visual distortions during image processing ($e.g.$, blur, compression, and noise) in the digital domain and environment variations ($e.g.$, lights, shooting angles, and distances) in the physical domain. It demonstrates that several advanced attacks with invisible triggers are sensitive to visual distortions since the trigger magnitude is small. In this case, a sufficiently large magnitude is desired to ensure that the trigger is not destroyed during the image processing, making a visible trigger more advantageous Additionally, the trigger added by the generative model can simulate distortions under different environmental conditions, thereby making the attack more robust.
@@ IV-B1 Trigger selection module
Coarse-grained selection. During the coarse-grained selection, triggers are filtered semantically to ensure compatibility within the image context. LLMs like GPT-4 [51] possess the capability to understand semantic information akin to humans. We first leverage their expansive prior knowledge to broadly filter out a list of candidate triggers. Specifically, for any given dataset, we provide the LLM with class names and prompt instructions, which outline the attacker’s semantic criteria for trigger selection, for example, “Find 10 common objects which look natural with $[classes]$”. Some examples are shown in supplementary material I-D.
@@ IV-B3 Quality assessment module
(a) ImageNet-Dogs (Target class: Maltese dog)
(b) FOOD-11 (Target class: Bread)
Figure 4: Poisoned samples generated by different attacks for the image classification task. BadNets uses a black-and-white grid as the trigger pattern, while Blended uses an image. For VSSC attack, we demonstrate two semantic triggers used in our experiments for each dataset. Although all these poisoned images are successfully classified as the target label, VSSC triggers in them are the most stealthy and compatible.
Remark on characteristics of the generated triggers. Several examples of poisoned images are illustrated in Figure 4. The semantic trigger has been effectively integrated into the benign image, with its shape, size, and location adjusted to ensure compatibility with the remaining visual content in the edited image. Consequently, we name the proposed trigger as a Visible, Semantic, Sample-specific, and Compatible trigger (VSSC trigger). As demonstrated in Table I, our trigger is the only one that satisfies visible, semantic, sample-specific, and compatible properties simultaneously, which aligns with the attack goals introduced in Section III. Furthermore, thanks to the automated trigger generation pipeline, VSSC attack can achieve an efficient physical attack without the need for heavy manpower.
@@ V-A Evaluation Tasks and Scenarios
  • Digital-to-physical (D2P) scenario: During the actual implementation of backdoor attacks in the physical domain, visual distortion may be unintentionally introduced. To evaluate the attack performance against such distortions, in this scenario, we simulate visual distortions by printing testing poisoned images and recapturing them, following the same way adopted in  [57].
@@ V-B1 Experimental settings
Dataset and models. We use two high resolution datasets: ImageNet-Dogs [58], a 20,250-image subset of ImageNet [59] featuring 15 dog breeds, and FOOD-11 [60], containing 16,643 images across 11 food categories. These two datasets are used to validate the effectiveness of VSSC triggers. The image size is 3$\times$224$\times$224. We use ResNet-18 [61] and VGG19-BN [62] for both datasets. The main results are based on ResNet-18, with more experiments with VGG19-BN in supplementary material II-A.
Implementation details. For the image classification task, the ISR threshold is set to 0.5 during the fine-grained selection. From the final selected triggers, we choose “red flower” and “harness” for the ImageNet-Dogs dataset and “nuts” and “strawberry” for the FOOD-11 dataset for experimental demonstration. The poisoning ratio is set to 5% and 10%, with a target label $y_{t}=0$ assigned to all datasets. Further details can be found in supplementary material I-A3.
@@
(a) Dogs with “red flower”
(b) Dogs with “harness”
(c) Food with “nuts”
(d) Food with “strawberry”
Figure 5: Examples of manually captured photos in the image classification task.
@@ V-B2 Effectiveness in digital scenario
As shown in Table II, our attack achieves an ASR up to 97.16% on the ImageNet-Dogs dataset and 97.22% on the FOOD-11 dataset with a poisoning rate of 5%. Although we use triggers with actual semantics and thus have representations closer to the original images in feature space, our attack method still demonstrates solid effectiveness in digital scenarios, comparable to other baseline attacks. It is worth emphasizing that VSSC attack has a minimal impact on C-Acc, causing less than 2% C-Acc decrease across all backbones and datasets, outperforming all other attacks. The sample-specific characteristic of VSSC triggers introduces a greater diversity in the inserted triggers, enabling the model to learn the common features of the triggers during training and to distinguish them from key features in normal image classification tasks. This makes the detection of VSSC attacks more challenging. Results with a ratio of 10% are in supplementary material II-A1.
@@ V-B3 Effectiveness in digital-to-physical scenario
As illustrated in Table III, VSSC attack achieves the highest ASR in the digital-to-physical scenario. On the ImageNet-Dogs dataset, the ResNet-18 model poisoned by VSSC triggers achieves an ASR of 97.62% in the digital-to-physical scenario with a poisoning ratio of 5%, and 93.33% on the FOOD-11 dataset. The VSSC trigger achieves the highest ASR among all attacks on both datasets. In contrast, none of the baseline attacks maintain effectiveness with an acceptable C-Acc, primarily because visual distortions are not accounted for during the design of their triggers, rendering them susceptible in practical deployments. In addition, VSSC attack has the lowest R-Acc on both datasets, which means that the backdoor model does not tend to classify poisoned samples into their original classes, also reflecting the effectiveness of the backdoor attack. Figure 14 presents the Grad-CAM [64] visualizations of images with the VSSC trigger and two visible triggers. When subjected to visual distortions, the attention regions of the backdoor model for both BadNets and TrojanNN shift, whereas the attention region associated with the VSSC trigger remains constant. This indicates that the VSSC trigger exhibits remarkable robustness against visual distortions.
@@ V-B4 Effectiveness in physical scenario
For experiments in the physical scenario, we collect 100 images with the trigger entity for each trigger. Figure 5 shows some examples of these manually captured images. Experiments are conducted under four poisoning ratios, ranging from 5% to 30%, as shown in Table IV. The VSSC trigger can implement successful attacks in the physical scenario, even at very low poisoning ratios. At a poisoning ratio of 5%, most triggers achieve an ASR of over 60%, which is notably high when utilizing a category of objects with specific semantics as triggers. With the increase in the poisoning ratio, the ASR correspondingly improves. At a poisoning ratio of 30%, the ASR of the “strawberry” trigger even reaches 100%. As a trigger deployed in the digital domain, the VSSC trigger still achieves a remarkably high attack success rate in the physical scenario, demonstrating its effectiveness in the real world.
Figure 6: Poisoned samples created by BadNets and VSSC for the object detection task. For the VSSC attack, we show “lemon” and “beachball” as triggers in three different scenarios, where D2P means the digital-to-physical scenario. We only use the results under Global Misclassification Attack (GMA) for demonstration.